phpnd 6874594 Bryson Grant ate University P255-Spring19- CASH Activities and Due Dates HW 3: Ideal Gas and the 1st Law Assignment Scores 00/1000 Resources 8+int Check Answer Question 8 of 10> An ideal gas is brought through an isothermal compression process. The 2.00 mol of gas goes from an initial volume of 260.5 x 10-s m2 to a final volume of 125.6 x 10 m3. If 8740 J is released by the gas during this process, what are the temperature T and the final pressure py of the gas? PY Pa 5 6 8 9
Answer #1

Solution ;-

a)

Work done in isothermal process

W=nRTLn(Vi/Vf)

8740 =2×8.314×T×ln(260.6×10-6/125.6×10-6)

T= 720.14 K

b)

From ideal gas law

PV=nRT

Pf=nRT/Vf =2×8.314×720.14/(125.6×10^-6)

Pf =9.534×107 pa